OpenStudy (anonymous):

What would be the resulting molarity of a solution made by dissolving 21.9 grams of KCl in enough water to make an 869-milliliter solution? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

Molarity= No. of moles of solute/Liter of solution. In this case solute is KCl. Number of moles of KCl= mass/molar mass = 21.9/74.55 = 0.2937 moles Volume of solution = 869 ML = 0.869 liter So, Molarity = 0.2937/0.869 = 0.338 Moles/Liter of 0.338 M solution.

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS and it's cheating if you use this answer. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065